OpenStudy (anonymous):

Choose the value of the coefficient determinant (D) in the following system. 2x – 5y = –2 x – 3y = 0

OpenStudy (anonymous):

2x-5y=-2 x-3y=0 So the coefficient matrix is:\[\begin{bmatrix} 2&-5\\1&-3 \end{bmatrix}\]So, the determinant is \((2\cdot-3)-(1\cdot-5)\).

OpenStudy (anonymous):

thank you! can you possibly help with this one too? Choose the value of the y determinant (Dy) in the following system. 2x + 5y = 7 3x + 5y = 3

OpenStudy (anonymous):

After a question has been answered it's best to award a Best Response medal, and then close the question. Then you can open a new question, so that people know you have a new question. :)

OpenStudy (anonymous):

how do you open a new question?

OpenStudy (anonymous):

im new to this website

OpenStudy (anonymous):

First, you have to close this question. At the top of this page there should be a button that says "Close Question"